OpenStudy (ckohler23):

What is the equation of a line passing through (-3,7) and having a slope of -1/5 /=fraction

OpenStudy (koikkara):

Can you please draw it, so that we can find answer together !

OpenStudy (rolldextro):

Y=32/5-x/5

MARC (marc_d):

Step 1: Use the formula\[\huge{y=mx+c}\]where =slope c=y-intercept using the information given\[(-3,7)=(x,y)\]and\[slope,m=-\frac{ 1 }{ 5 }\]substitute those value into the formula\[y=mx+c\]to find the value of c(y-intercept)

MARC (marc_d):

\[\huge{7=-\frac{ 1 }{ 5 }(-3)+c}\]remember that \[(-)negative~value~\times(-)negative~value=(+)postive~value\]\[(+)positive~value~\times(+)positive~value=(+)postive~value\]\[(-)negative~value~\times(+)positive~value=(-)negative~value\]\[(+)positive~value~\times(-)negative~value=(-)negative~value\]

MARC (marc_d):

\[\huge{7=\color{red}{-}\frac{ 1 }{ 5 }(\color{red}{-}3)+c}\]\[\huge{7=\frac{ 3 }{ 5 }+c}\]Then,subtract both sides by 3/5\[\huge{7-\frac{ 3 }{ 5 }=\frac{ 3 }{ 5 }-\frac{ 3 }{ 5 }+c}\]\[\huge{c=6\frac{ 2 }{ 5 }}\]

MARC (marc_d):

\[\huge{c=\frac{ 32 }{ 5 }}\]

MARC (marc_d):

Step 2: Substitute the value of slope,m and c(y-intercept)into the formula\[\huge{y=mx+c}\]to form the equation

MARC (marc_d):

@Ckohler23